# Profile-Aware Maneuvering: A Dynamic Multi-Agent System for Robust GAIA Problem Solving by AWorld

### v3

@@ Abstract
The rapid advancement of large language models (LLMs) has empowered intelligent agents to leverage diverse external tools for solving complex real-world problems. However, this reliance introduces new challenges, as extended contexts and noisy tool outputs can undermine system reliability. To address this, we propose a dynamic Multi-Agent System (MAS) in our AWorld framework, where an Execution Agent is supervised by a Guard Agent that provides on-demand "dynamic maneuvering"—verifying and correcting the reasoning process to improve robustness over single-agent systems. To move beyond this generic supervision, we enhance the architecture with a methodology inspired by System Identification from control theory. This method first profiles the Execution Agent offline on a benchmark dataset to create a "performance fingerprint" of its unique weaknesses. The Guard Agent then leverages this fingerprint online to deliver profile-aware supervision, making targeted interventions based on known failure patterns rather than merely reacting to immediate logical flaws. Extensive experiments on the GAIA dataset demonstrate that this profile-aware MAS significantly improves both effectiveness and stability, outperforming not only single-agent systems but also its ’naive’ counterpart. This superior performance led our system to achieve first place among open-source projects on the prestigious GAIA leaderboard. These findings highlight that building truly trustworthy intelligent systems requires not just collaboration, but a deep, empirically-grounded understanding of each agent’s unique capabilities and limitations.
@@ 1 Introduction
However, to elevate this system from merely reactive correction to proactive, intelligent guidance, we address its key limitation: a generic Guard Agent, while beneficial, remains blind to its partner’s specific, habitual failure modes and cannot thus provide more targeted supervision and guidance. We overcome this by introducing our core contribution, inspired by System Identification—a mature engineering discipline effectively applied in domains like vessel maneuvering research where understanding dynamic behavior is paramount (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013). Our adaptation of this philosophy to model LLM agents bifurcates into two distinct phases. First, in a preparatory offline stage, we systematically benchmark the Execution Agent to generate a detailed "performance fingerprint", a structured profile of its characteristic errors, such as a tendency to hallucinate code. Then, in the crucial online execution stage, the Guard Agent leverages this fingerprint to provide profile-aware maneuvering, actively monitoring for likely failure scenarios and offering targeted, preemptive advice. This transforms the Guard Agent into an expert on its specific partner.
@@ 2 Method
2.1 The Multi-Agent System and Dynamic Maneuvering
Our approach is inspired by well-established principles in control theory, particularly from the field of marine vessel navigation (Xie et al., 2020). The maneuvering of a ship, as modeled by the control equations in Table 1, represents a system subject to complex external forces like wave drift. To ensure the vessel converges to a desired trajectory (e.g., the zig-zag test in Figure 3), continuous, adaptive adjustments are essential. The rudder angle is dynamically controlled not with a fixed setting, but to actively counteract disturbances and guide the ship’s path.
We posit that an agent’s problem-solving process is analogous to this voyage. The agent’s reasoning forms a "trajectory" toward a solution, constantly perturbed by "noise" from long contexts and irrelevant tool outputs—akin to the winds and currents affecting a vessel. This analogy directly informs our foundational system design within the AWorld framework, where we implement a dynamic maneuvering mechanism. As shown in Figure 4, we establish a collaborative dyad:
An Execution Agent, which performs the primary task.
  • An Execution Agent, which performs the primary task.
  • A Guard Agent, which acts as a dynamic supervisor, analogous to a ship’s rudder.
Just as a rudder provides corrective steering, the Execution Agent can invoke the Guard Agent at critical junctures to verify logic and correct deviations, ensuring it remains on a robust path to the final goal.
$(X_{\dot{u}}-m)\dot{u}+X_{u}\Delta u+X_{\text{drift}}=0$ (1)
$(Y_{\dot{v}}-m)\dot{v}+Y_{v}v+(Y_{\dot{r}}-mx_{G})\dot{r}+(Y_{r}-mu_{0})r+Y_{\text{drift}}=0$
$(N_{\dot{v}}-mx_{G})\dot{v}+N_{v}v+(N_{\dot{r}}-I_{Z})\dot{r}+(N_{r}-mx_{G}u_{0})r+N_{\text{drift}}=0$
Table 1: In these linearized equations of motion, terms like $u,v,r$ represent velocities. The key insight for our work lies with the hydrodynamic coefficients (e.g., $X_{\dot{u}}$, $Y_{v}$, $N_{r}$). These coefficients are not universal; they are unique to each ship’s specific hull geometry and mass distribution and must be determined experimentally or numerically. This process of identifying a system’s unique internal parameters from observed behavior is the essence of System Identification. (Alexandersson et al., 2024; Xue et al., 2021; Xu and Soares, 2013)
Figure 3: Phase portrait where the combination of drift angle and yaw rate is shown for zigzag10/10 and zigzag20/20 wPCC model tests. (Alexandersson et al., 2024)
2.2 From Generic Maneuvering to Profile-Aware Guidance
While the foundational MAS architecture improves robustness, its supervision is generic. The Guard Agent reacts to logical errors as they arise but is blind to the Execution Agent’s habitual weaknesses. To achieve a deeper level of intelligence, we draw a more profound inspiration from the practice of System Identification.
2.2.1 Agent Profiling via System Identification
In naval architecture, the equations in Table 1 provide a general model of ship motion. However, to be useful, the specific values of the hydrodynamic coefficients ($X_{\dot{u}},Y_{v}$, etc.) must be determined for each individual vessel. These coefficients, unique to each ship’s hull and design, are identified through a series of standardized experimental tests (like the zig-zag test). This process creates a precise mathematical "fingerprint" of that ship’s unique maneuvering behavior.
We contend that LLM-based agents, like ships, are complex systems with unique, idiosyncratic behaviors. A foundational model provides the general architecture, but how a specific agent instance—shaped by its system prompt and tools—responds to different situations is analogous to a ship’s unknown hydrodynamic coefficients. We therefore adapt the philosophy of System Identification to create an agent profiling methodology as shown in Figure 4. This involves two steps:
  • Systematic Benchmarking: In a preparatory offline stage, we subject the Execution Agent to 62 problems from the GAIA validation dataset (similar distribution but distinct from the test dataset). This is analogous to putting a ship through its sea trials.
  • Fingerprint Generation: We then task a separate, high capacity LLM-based analyzer (Gemini 2.5 Pro in our practice) to study the Execution Agent’s complete input-output logs from the benchmark. The analyzer’s goal is not merely to score performance but to identify and articulate characteristic failure modes. The identification system prompt can be referred in the appendix. It synthesizes these observations into a structured, human readable profile of an agent’s weaknesses—its "performance fingerprint." This fingerprint might detail a tendency to "hallucinate code when a web search fails," a "proclivity for off-by-one errors in array manipulation," or an "over-reliance on the first search result, even if it’s from an unreliable source."
Figure 4: Multi-Agent System built on the Aworld framework: There are three main configurations evaluated in our experiments that can be derived from this figure. Single Agent System: The baseline system, where an Execution Agent uses a set of tools without the Guard Agent. Naive MAS: A Guard Agent is introduced to provide reactive, real time logical verification for the Execution Agent. Profile-Aware MAS: The final system, where the Guard Agent is equipped with a performance fingerprint derived from the System Identification through the offline process, enabling proactive and targeted supervision.
2.2.2 Profile-Aware Supervision in Online Execution
The generated performance fingerprint is the key to transforming the Guard Agent from a naive supervisor into an expert partner. During online task execution, the Guard Agent is "armed" with this fingerprint, typically by injecting it into its system prompt as a core piece of contextual knowledge, as shown in Figure 4.
This fundamentally alters its behavior. Instead of performing generic logical checks, the Guard Agent now provides profile-aware supervision. It actively monitors for scenarios where the Execution Agent is statistically likely to fail according to its profile. For instance, if the fingerprint identifies a weakness in handling multi-step calculations, the Guard Agent will apply heightened scrutiny and may proactively request a step-by-step breakdown whenever such a calculation is initiated. This shift turns its role from reactive correction to targeted, preemptive guidance, directly addressing the agent’s known blind spots and dramatically enhancing the reliability of the entire system.
@@ 4 Experimental Results
Round 1 Pass@1 32.11% 56.88% 70.64% 72.48%
Round 2 Pass@1 30.28% 63.30% 64.22% 70.64%
Round 3 Pass@1 32.11% 64.22% 66.06% 69.72%
@@
Table 2 demonstrates the experimental results, showing a progressive improvement across the four methodologies.
The Base model (Gemini 2.5 Pro), which, relying solely on internal knowledge, achieves an average pass@1 accuracy of 31.50%. Introducing tools in the Single Agent System (SAS) nearly doubles this figure, raising the pass@1 to 61.47% by enabling context expansion and real-world data acquisition. Building on this, our Multi-Agent System (MAS) leverages a "naive" Guard Agent for real-time verification, further increasing the pass@1 accuracy to 66.97%, an 8.95% improvement over the SAS. Finally, equipping the Guard Agent with the performance fingerprint in our Profile-Aware MAS pushes these accuracy gains to their peak. This final enhancement further elevates the average pass@1 to 70.95% and the pass@3 to 84.40%. This demonstrates that profile-aware supervision, which proactively targets known failure modes, provides a distinct and measurable advantage over the reactive, generic checks of the naive MAS.
Beyond accuracy, the results reveal the improvement about system stability. While the Base model is highly consistent (standard deviation of 0.00863), the SAS experiences a nearly fourfold increase in score variance, a direct consequence of the uncertainties introduced by external tools. The naive MAS begins to counteract this instability, with the Guard Agent’s logical constraints reducing the standard deviation by 17.18% compared to the SAS. This stabilizing effect becomes dramatically more pronounced with profile-aware maneuvering. The Profile-Aware MAS slashes the standard deviation by an additional 57.41% down to 0.01147, bringing the system’s consistency nearly back to the level of the Base model but at a vastly superior accuracy. This highlights the crucial difference between merely correcting errors after they occur and proactively preventing them based on an empirical understanding of the agent’s behavior.

### v4

@@ Abstract
The rapid advancement of large language models (LLMs) has empowered intelligent agents to leverage diverse external tools for solving complex real-world problems. However, this reliance introduces new challenges, as extended contexts and noisy tool outputs can undermine system reliability. To address this, we reframe agent design as a control engineering discipline. We first introduce a baseline Multi-Agent System (MAS) that acts as a simple reactive feedback controller, where a Guard Agent corrects the primary Execution Agent’s errors. Furthermore, the core of our contribution is a more sophisticated architecture inspired by System Identification. We develop an automated offline process to generate a "performance fingerprint"—an explicit model of the Execution Agent’s characteristic failure modes. Armed with this fingerprint, the Guard Agent evolves into a predictive controller, implementing a feed-forward strategy to preemptively counteract errors before they derail the reasoning process. Experiments on the challenging GAIA benchmark validate our hierarchical approach. The final Profile-Aware MAS demonstrates the hallmarks of a well-controlled system: it dramatically reduces performance variance and minimizes the gap between its potential and single-pass performance, enhancing overall reliability. This superior performance and stability led our system by AWorld to achieve first place among open-source projects on the GAIA leaderboard. Our findings advocate for a paradigm shift: from the empirical craft of prompt engineering to the rigorous discipline of control theory for designing predictable and trustworthy intelligent agents.
@@ 1 Introduction
However, to elevate this system from merely reactive correction to proactive, intelligent guidance, we address its key limitation: a generic Guard Agent, while beneficial, remains blind to its partner’s specific, habitual failure modes and cannot thus provide more targeted supervision and guidance. We overcome this by introducing our core contribution, inspired by System Identification—a mature engineering discipline effectively applied in domains like vessel maneuvering research where understanding dynamic behavior is paramount (Xu and Soares, 2013; Xue et al., 2021; Alexandersson et al., 2024). Our adaptation of this philosophy to model LLM agents bifurcates into two distinct phases. First, in a preparatory offline stage, we systematically benchmark the Execution Agent to generate a detailed "performance fingerprint", a structured profile of its characteristic errors, such as a tendency to hallucinate code. Then, in the crucial online execution stage, the Guard Agent leverages this fingerprint to provide profile-aware maneuvering, actively monitoring for likely failure scenarios and offering targeted, preemptive advice. This transforms the Guard Agent into an expert on its specific partner.
@@ 2 Method
Our methodology is grounded in the principles of Engineering Cybernetics and Control Theory, which provide a robust framework for designing and analyzing complex, dynamic systems. We first establish our theoretical foundation by drawing a parallel with the well-understood problem of marine vessel maneuvering. We then formalize the LLM-based agent as a controllable system. Finally, we detail our progressively sophisticated control architectures, interpreting them through the rigorous lens of control theory.
2.1 Theoretical Foundation: Maneuvering Complex Systems
The core challenge in controlling any complex system, from a supertanker to an intelligent agent, is to ensure its behavior converges to a desired trajectory despite internal instabilities and external disturbances. In Engineering Cybernetics, this is achieved by first understanding the system’s intrinsic dynamics and then designing an appropriate control strategy (Tsien and Qian, 1954).
A perfect, tangible illustration of this principle is found in marine vessel navigation (Xie et al., 2020; Alexandersson et al., 2024). The motion of a ship is governed by a set of linearized equations. Critically, these equations contain hydrodynamic coefficients ($X_{\dot{u}},Y_{v}$, etc.) that are unique to each vessel. These unknown parameters must be experimentally determined through a process called System Identification, where standardized tests (like the zig-zag maneuver shown in Figure 3) are used to create a precise mathematical "fingerprint" of the ship’s behavior. Only with this fingerprint can an effective control system (e.g., an autopilot) be designed to actively counteract disturbances and guide the ship along a desired path.
We posit that this two-stage philosophy—first, identify the system’s unique behavioral fingerprint, then design a targeted control architecture—is directly applicable to engineering reliable LLM-based agents.
Figure 3: The zig-zag test is a standard procedure in System Identification for marine vessels, designed to reveal the ship’s unique maneuvering characteristics (its "fingerprint") (Alexandersson et al., 2024).
2.2 Modeling the Agent as a Controllable System
To apply the rigorous discipline of control theory, we first formalize the agent’s problem-solving process as a controllable system. The foundational LLM, a static function $y=f_{\theta}(x)$ with immutable parameters $\theta$, serves as the core of our system. Our entire methodology is an exercise in applied control engineering: how to strategically design the input context, $x$, to steer the output, $y$, towards a correct and stable solution.
Before detailing our architectures, we establish a mapping between control theory variables and their concrete instantiations within our agent framework:
  • Plant ($P$): The Execution Agent ("E"), the core process we aim to control. Its behavior is dictated by $f_{\theta}$.
  • Controlled Variable ($y$): The Execution Agent’s output, $y_{E}$. This is the variable we want to regulate.
  • Setpoint ($r$): The implicit, desired "correct" reasoning path. This is the target trajectory.
  • Error ($e$): The deviation of the agent’s output from the correct path, $e=r-y_{E}$. This is often detected implicitly as logical fallacies or factual inaccuracies.
  • Disturbance ($d$): Internal factors (e.g., hallucinations, logical fallacies) and external factors (e.g., noisy tool outputs) that cause $y_{E}$ to deviate from $r$.
  • Controller ($C$): The Guard Agent ("G"), which observes the system and computes a corrective action. Its behavior is also dictated by $f_{\theta}$.
  • Control Signal ($u$): The critique or guidance $y_{G}$ generated by the Guard Agent. This is the action applied to steer the plant.
2.3 Hierarchical Control Architectures for Agent Maneuvering
We designed and evaluated a hierarchy of control architectures, each representing a more sophisticated control strategy, as depicted in Figure 4.
Figure 4: Our hierarchical control architectures, built on the Aworld framework. This figure illustrates the components for the Single Agent System (uncontrolled), the Naive MAS (feedback control), and the Profile-Aware MAS (composite feed-forward-feedback control), which leverages a fingerprint from an offline System Identification process.
2.3.1 The Uncontrolled System (Single Agent System)
In the baseline case, the agent operates in an "open loop" without a controller. The control signal is effectively zero ($u=0$). At each step $t$, the agent’s context $x_{t}$ is formed by concatenating the task $x_{\text{task}}$, its prior reasoning $\mathcal{T}_{t}=\{y_{E,0},...,y_{E,t-1}\}$, and new tool information $\text{Info}_{t}$. The agent’s next action is then:
$y_{E,t}=f_{\theta}(x_{E,t})\quad\text{where}\quad x_{E,t}=x_{\text{task}}\oplus\mathcal{T}_{t}\oplus\text{Info}_{t}$
This system relies solely on the intrinsic capabilities of the plant ($f_{\theta}$) and is highly susceptible to disturbances, analogous to a ship drifting without rudder control.
2.3.2 Reactive Feedback Control (The Naive MAS)
Our first control strategy implements a classic negative feedback loop. A Guard Agent ("G") is introduced to act as the controller. In standard control theory, the controller’s action is defined by the feedback law $u=C(e)$. In our framework, this is realized as follows:
  • The Guard Agent observes the plant’s output trajectory, $\mathcal{T}_{E,t+1}$.
  • By analyzing this trajectory for logical flaws, it implicitly computes the error $e$.
  • It then generates a critique, which serves as the control signal $u$.
This process is formally described as:
$u_{t}=y_{G,t}\quad\text{where}\quad y_{G,t}=f_{\theta}(\mathcal{T}_{E,t+1})$
This control signal $u_{t}$ is then fed back into the Execution Agent’s next context, creating the closed loop:
$x_{E,t+1}=x_{E,t}\oplus u_{t}$
This architecture mirrors a standard feedback controller, which is effective for stabilization but is fundamentally reactive, it can only correct an error after it has already occurred and been detected.
2.3.3 Predictive Composite Control (The Profile-Aware MAS)
The pinnacle of our approach is a sophisticated composite feed-forward-feedback control system. This strategy first requires creating a model of the plant’s predictable behaviors via System Identification (Xu and Soares, 2013; Xue et al., 2021; Alexandersson et al., 2024).
  • Offline System Identification: We subject the Execution Agent ("P") to benchmark problems (GAIA validation dataset), analyzing its logs to identify characteristic failure modes in response to certain types of tasks (disturbances, "d").
  • Generating the Fingerprint: This analysis is synthesized into a performance fingerprint, $\mathcal{P}_{E}$.
In the online phase, this fingerprint enables a composite control law, $u=u_{fb}+u_{ff}$, where $u_{fb}$ is the reactive feedback signal and $u_{ff}$ is a proactive feed-forward signal. Our non-linear controller $f_{\theta}$ generates this composite signal in a single inference step:
$u_{t}^{\prime}=y^{\prime}_{G,t}\quad\text{where}\quad y^{\prime}_{G,t}=f_{\theta}(\underbrace{\mathcal{T}_{E,t+1}}_{\text{for }u_{fb}}\oplus\underbrace{\mathcal{P}_{E}}_{\text{for }u_{ff}})$
Here, the Guard Agent processes $\mathcal{T}_{E,t+1}$ to generate the reactive feedback component ($u_{fb}$) for any observed error, while simultaneously processing $\mathcal{P}_{E}$ to anticipate errors characteristic of the Execution Agent and generate a pre-emptive feed-forward component ($u_{ff}$). This proactive guidance is the hallmark of a truly robust control system.
2.3.4 Theoretical Justification of Composite Control
We now provide a formal justification from control theory to prove why a composite system is theoretically superior for handling predictable disturbances. Using Laplace transforms, where functions of time $t$ become functions of a complex variable $s$, we analyze the system’s response to a disturbance $D(s)$.
Limitation of Pure Feedback Control (Naive MAS):
In a standard feedback system, the relationships between the components are defined as follows: the system output $Y(s)$ is the sum of the plant’s action and the disturbance; the control signal $U(s)$ is the controller’s action on the error; and the error $E(s)$ is the difference between the setpoint $R(s)$ and the output.
$Y(s)$ $=P(s)U(s)+D(s)$ (1)
$U(s)$ $=C(s)E(s)$ (2)
$E(s)$ $=R(s)-Y(s)$ (3)
By substituting (3) into (2), and then into (1), we can solve for the closed-loop output $Y(s)$:
$Y(s)=\frac{P(s)C(s)}{1+P(s)C(s)}R(s)+\frac{1}{1+P(s)C(s)}D(s)$
From this equation, it is evident that the system can only suppress the disturbance $D(s)$ by increasing the controller gain to make the term $1+P(s)C(s)$ large. It cannot eliminate the disturbance’s effect entirely, as the controller only acts after $D(s)$ has already corrupted the output $Y(s)$. This is the mathematical definition of a reactive system.
Superiority of Composite Control (Profile-Aware MAS):
With a feed-forward controller $C_{ff}$ added, the total control signal becomes a composite of the feedback signal $U_{fb}(s)$ and the feed-forward signal $U_{ff}(s)$.
$U(s)$ $=U_{fb}(s)+U_{ff}(s)$ (4)
$U_{fb}(s)$ $=C_{fb}(s)E(s)=C_{fb}(s)(R(s)-Y(s))$ (5)
$U_{ff}(s)$ $=C_{ff}(s)D(s)$ (6)
Substituting this composite control law into the plant equation (1):
$Y(s)=\frac{P(s)C_{fb}(s)}{1+P(s)C_{fb}(s)}R(s)+\frac{1+P(s)C_{ff}(s)}{1+P(s)C_{fb}(s)}D(s)$
This equation reveals a theoretical breakthrough. Perfect rejection of the disturbance is theoretically possible if we can design a feed-forward controller $C_{ff}(s)$ such that the numerator of the disturbance term becomes zero:
$\text{If }C_{ff}(s)=-\frac{1}{P(s)},\quad\text{then the term }1+P(s)C_{ff}(s)=0$
This would cancel the disturbance preemptively, before it affects the output.
The performance fingerprint $\mathcal{P}_{E}$ serves as our empirically-derived model of the plant $P(s)$. The ideal feed-forward controller, $C_{ff}(s)=-1/P(s)$, requires a controller that acts as the negative inverse of the plant. In our system, this "negative inverse" is not achieved via a literal sign but through the learned function of the Guard Agent. The Guard Agent’s role is defined as a critiquer and corrector. Its fundamental objective is to generate an output, a critique, whose semantic effect is to invert and cancel the predicted error of the Execution Agent. This core corrective purpose of the Guard Agent embodies the negative sign required by control theory, ensuring the feed-forward action suppresses, rather than amplifies, disturbances. This formal analysis proves that our profile-aware architecture is theoretically superior for achieving robust, predictive, and stable agent maneuvering.
@@ 4 Experimental Results
Round 1 P@1 32.11% 56.88% 70.64% 72.48%
Round 2 P@1 30.28% 63.30% 64.22% 70.64%
Round 3 P@1 32.11% 64.22% 66.06% 69.72%
@@
The experimental results, summarized in Table 1, quantitatively validate our control-theoretic approach, demonstrating a clear, progressive improvement across three key dimensions: raw accuracy, system stability, and reasoning reliability.
Accuracy Progression: A Stepwise Ascent. The data shows a clear, stepwise improvement in performance with each architectural enhancement. The Base model (Gemini 2.5 Pro), relying solely on its parametric knowledge, establishes a baseline Pass@1_avg of 31.50%. Introducing tools in the Single Agent System (SAS) delivers the largest single leap, nearly doubling the accuracy to 61.47% by enabling real-world data acquisition. Building on this, the Multi-Agent System (MAS), through its reactive feedback control, elevates the Pass@1_avg to 66.97%—an 8.95% relative improvement. Finally, the Profile-Aware MAS (PA-MAS), leveraging predictive feed-forward control, pushes the accuracy to its peak of 70.95%. The consistent climb towards a final Pass@3 of 84.40% confirms that each layer of our control architecture contributes distinct and measurable value.
System Stability: Taming the Chaos of Capability. Beyond raw accuracy, the standard deviation (Pass@1_std) tells a compelling story about stability. The SAS, while more capable, pays a steep price: its score variance explodes by nearly 280% compared to the Base model. This reflects the classic engineering trade-off where adding capabilities (tools) introduces significant uncertainty and instability. The MAS architecture marks the turning point, where the Guard Agent’s feedback begins to rein in this chaos, reducing the standard deviation by 17.18%. This stabilizing effect becomes dramatically more pronounced with our final architecture. The PA-MAS slashes the standard deviation by a further 57.41%, bringing the system’s consistency nearly back to the level of the hyper-stable Base model, but at a vastly superior performance level. This validates that our profile-aware control architecture is not just correcting errors, but fundamentally stabilizing the agent’s behavior.
Reasoning Reliability: Closing the Gap Between Potential and Performance. Perhaps the most profound insight comes from analyzing the gap between a system’s potential (Pass@3) and its average first-shot success (Pass@1_avg). This ‘Pass@3 - Pass@1_avg‘ metric quantifies the "regret" of a system—how often it fails on a single attempt despite possessing the underlying capability to succeed.
  • The SAS exhibits the largest gap at 19.26% (80.73% - 61.47%). It has high potential but is highly unreliable, frequently failing to realize that potential on the first try.
  • The Naive MAS, with its generic feedback, narrows this gap to 15.60% (82.57% - 66.97%), improving first-shot reliability.
  • The Profile-Aware MAS achieves the tightest gap of all agentic systems at 13.45% (84.40% - 70.95%).
This demonstrates that the Profile-Aware MAS is not only the most accurate and stable system but also the most reliable in its reasoning. It maximizes its performance potential in a single pass, minimizing failures due to stochasticity or predictable error paths. This trifecta of improvements—higher accuracy, lower variance, and a smaller potential-performance gap—provides conclusive evidence that applying a profile-aware, composite control strategy is a superior paradigm for engineering dependable intelligent agents.